# Sea ice floe segmentation in close-range optical imagery using active contour and foundation models

### v3

@@ 4.3 A combination of SAM and GVF (SAM-GVF)
Our analysis indicates that the GVF algorithm efficiently delineates floes but encounters difficulties with varied gradients in gray-scale intensity due to its reliance on visual characteristics. SAM, on the other hand, can process a wide range of image variations due to its extended training, but it lacks shape integration, making it difficult to segment adjacent floes with weak edges. We propose combining SAM with GVF to compensate for shortcomings of the former by incorporating shape details, an approach conceptually similar to those suggested previously (Chen et al., 2019; Hoang Ngan Le et al., 2020, e.g.,).
@@ 5.1 Benchmark dataset
A benchmark is created by manually outlining a subset of twenty images from the original dataset (see §B), which span a wide range of visual conditions and varying levels of sea ice clarity. Segmentation of individual floes is performed using image editing software (Pixelmator, though other tools could also be used). Reference data for the number and size of floes, as well as ice concentration, are subsequently derived. Concurrently, the same sample is segmented using the automated segmentation methods (GVF, SAM-A, SAM-P, and the combined SAM-GVF). Over 2,800 floes were identified across the 20 images, providing a solid foundation for benchmarking (Arbelàez et al., 2011; Martin et al., 2001).
@@ 5.2 Visual analysis of segmentation performance on sample images
Two sample images representing contrasting ice morphology, one with distinct ice floes and well-defined edges, and another featuring diffuse boundaries due to closely packed or interconnected floes and interstitial ice, are used to assess the different methods against manual segmentation. This comparison illustrates how varying sea ice conditions influence segmentation outcomes and provides insight into model performances (within their respective constraints). These include: missing floes, where floes identified in the manual segmentation are not detected by the automated methods; misinterpreted floes, which are either partially segmented or merged with adjacent floes; and “hallucinations”, referring to false objects where patterns or textures are mistakenly identified as floes.
@@ 5.2.1 Distinct ice floes with defined edges
SAM-GVF misses ten floes, misinterprets four, and produces seven hallucinations ($\approx 10\%$ of floes). The missing floes have an average diameter of 2.6 m, covering around 1% of the surface area, similar to SAM-A and SAM-P. Misinterpreted floes have an average diameter of 3.4 m and include both partially segmented and merged floes. Hallucinations have an average diameter of 1.8 m and align with those seen in SAM-P.
Overall, SAM-GVF has the lowest percentage of segmentation errors, with uncertainties affecting $<$3% of the image area. SAM-P achieves similar performance ($\approx$6%), while SAM-A and GVF exhibit the highest uncertainty, affecting 11% and 12% of the surface area, respectively.
@@ 5.2.2 Diffuse sea ice boundaries
As expected, segmentation performance declines under more challenging conditions. For the sample image with diffuse boundaries, the manual segmentation detects 250 floes with an average diameter of 3 m and an ice concentration of 55%. The performance of GVF significantly degrades—it misses 104 floes, which is more than twice the number observed in the distinct edges image. Misinterpretations also increase, with 40 floes affected, while the 17 hallucinations is similar to the distinct edges case (Fig.9g). In total, $\approx$70of floes are not interpreted correctly, where the missed floes have an average diameter of 1.8 m and cover approximately 7% of the image area. The misinterpretations have an average diameter of 3 m, and result mostly from partial segmentation of larger floes, leading to a slight bias in favor of larger floes. The hallucinations have an average diameter of 1.1 m and do not substantially affect the overall sea ice characteristics. However, the increased number of missed and misinterpreted floes reduces the reliability of ice concentration estimates in the diffuse boundaries image.
@@
SAM-A is more robust for the diffuse boundaries image in comparison to GVF, although it shows a significant decline in performance. Approximately 81 floes are missed, which is over four times more than for the distinct edges image. Misinterpretations remain relatively stable at twenty floes, while hallucinations increase to ten, although this is still fewer than GVF (Fig.8e). Overall, around 45% of floes are not captured correctly. The missed floes have an average diameter of 1.3 m, covering about 3% of the image area. While SAM-A misses more floes in the diffuse boundaries image compared to the distinct edges image, it still retains the capacity to provide an accurate estimate of ice concentration. Misinterpretations have an average floe diameter of 4.5 m, and primarily arise from the merging of adjacent floes, continuing SAM-A’s tendency to bias floe size distribution toward larger floes. Despite the increased hallucinations, their overall impact on sea ice characteristics remains limited.
@@
SAM-P also experiences an increase in segmentation errors, missing 53 floes, which is greater than five times more than in the distinct edges sample. The missed floes are small, with an average diameter of 1.2 m and covering only 2% of the total area. Misinterpretations double to 22 instances, involving smaller floes (average diameter of 3.3 m) than those in the distinct edges image. These misinterpretations primarily result from partial segmentation of floes featuring both snow-covered and bare ice surfaces, which subsequently affects the estimated floe size. Hallucinations increase slightly to 16, likely due to the model’s sensitivity to complex textures when using prompts to refine uncertain regions. Approximately 36% of floes are not interpreted correctly.
@@
SAM-GVF shows moderate performance degradation, missing 58 floes, which is a slight increase compared to the distinct edges image. Misinterpretations remain relatively low, with 20 floes affected, while hallucinations increase slightly to 12, which is fewer than SAM-A and GVF (Fig.9g). Similar to SAM-P, $\approx$36of the floes are not captured correctly. The missed floes have an average diameter of 1.2 m, covering about 2% of the image area. Misinterpretations, with an average floe diameter of 2.8 m, mainly result from partial segmentation of larger floes and those with mixed snow and bare ice surfaces, which slightly distorts the floe size distribution. Hallucinations remain small, with an average diameter of 1.5 m, and do not significantly impact the overall sea ice characteristics.
@@
Of the various methods, SAM-P performs best across both the distinct edges and diffuse boundaries images, exhibiting fewer missed floes and relatively stable rates of misinterpretations and hallucinations. SAM-GVF also demonstrates good performance, while SAM-A and GVF show significant deterioration, particularly with a substantial increase in missed floes. Additionally, computational time per floe increases for GVF (and consequently for SAM-GVF) in the diffuse boundaries image, whereas it remains steady in SAM.
@@ 5.3 Floe count, size distribution, and sea ice concentration across the entire sample
Figure 10: Comparison of floe size statistics derived from segmentation methods GVF (green), SAM-A (red), SAM-P (pink), and SAM-GVF (purple) against the manual benchmark (orange). (a) Relative error (RE) in the number of detected floes compared to the manual benchmark (horizontal orange line). b) Violin plots of floe size distributions for each method, with the manual benchmark on the left. Squares inside violins represent medians; darker shaded areas indicate quartiles. (c–f) QQ plots comparing deciles (10% to 90%) of floe size distributions from segmentation methods against the manual benchmark. The solid line indicates perfect alignment. (g) Relative error (RE) in ice concentration compared to the manual benchmark (horizontal orange line).
@@
Data extracted from all 20 images in the sample are compared with the benchmark manual segmentation, as illustrated in Fig. 10. The relative error (RE) in estimating floe numbers is presented in Fig.10a. Among the methods evaluated, SAM-P performs the best, slightly underestimating the number of floes by approximately 2%, which highlights its capability to detect floes across images with varying clarity. In contrast, SAM-A and GVF are the least accurate, underestimating the manual count by approximately 28% and 26%, respectively. This reflects the difficulty in detecting floes in complex scenarios, particularly when dealing with unclear or irregular ice formations (cf. Figs. 8 and 9). For SAM-A, the automated initialization is a key limitation, as it prioritizes larger, more distinct floes while often overlooking smaller or poorly defined ones. The combined SAM-GVF approach mitigates some of the uncertainties in floe detection observed in both SAM-A and GVF, leading to a lower underestimation of the manual count (approximately 12%). This aligns with its performance on the distinct edges image, suggesting that, like SAM-P, it remains relatively consistent across varying image conditions.
@@
The floe size distribution is shown in Fig. 10b as violin plots, which display the overall distribution of data along with summary statistics, such as the median (marked inside the violins) and the first and third quartiles (represented by the darker shaded areas). The benchmark size distribution from manual segmentation exhibits a clear bimodal shape, with peaks around diameters of 1.5 m and 3.5 m. Approximately 50% of the floes fall between 1.8 m and 4.3 m, with a median diameter of 3 m. SAM-P aligns well with the benchmark, with the bimodal shape, median, and quartiles generally well captured. However, the distribution is slightly skewed toward smaller floes, as the first quartile underestimates the benchmark (1.5 m instead of 1.8 m), exaggerating the presence of smaller floes and reflecting the relatively large count of hallucinations (Fig. 8g). Both GVF and SAM-GVF methods also capture the overall shape of the distribution, including the median (3.2 m and 3.1 m, respectively) and quartiles. Minor discrepancies arise from the lower detection of smaller floes, resulting in a less pronounced bimodal shape compared to SAM-P. Conversely, SAM-A produces the least accurate distribution, missing the bimodal nature of the benchmark, resulting in a single peak around 3.3 m. It overestimates both the median (3.8 m) and quartiles, reflecting the limitations of SAM-A in detecting individual floes without seed-based prompts.
@@
The comparison between benchmark and model distributions is further illustrated and analyzed in Figs.10d-g, which present the deciles of the distributions as QQ plots. Despite differences in the overall bimodal shape, the deciles of SAM-P, SAM-GVF, and GVF align reasonably well with the benchmark (Figs.10c,e,f). The lowest deciles (10–30%) are slightly more pronounced, especially for SAM-P, suggesting that these methods detect a large number of small floes, attributed to spurious features or noise. However, the higher deciles align well with the benchmark. The root mean squared error (RMSE) of floe diameter is 0.20 m for SAM-P, 0.17 m for SAM-GVF, and 0.21 m for GVF. Given that the pixel-to-meter conversion in the orthorectification process was set to 0.05 m, the RMSE is approximately 3–4 times the smallest measurable unit in the images. SAM-A (Fig. 10d) shows a more significant underestimation compared to the other models, especially in the lowest five deciles. The RMSE for SAM-A is 0.94 m, reflecting its misalignment with the benchmark’s bimodality.
@@ 6 Conclusion
Three segmentation approaches have been evaluated for analyzing close-scale optical imagery of sea ice. The first approach employs the gradient vector flow (GVF) snake algorithm, which uses edge-based energy minimization for contour detection. The second approach utilizes the Segment Anything Model (SAM), a deep learning framework for image segmentation, applied in both its automatic mode (SAM-A) and a prompt-guided approach (SAM-P). The third approach combines SAM-P with the GVF algorithm (SAM-GVF) to leverage their complementary strengths. To ensure a robust and systematic evaluation, we implemented a fully automated image processing pipeline, incorporating orthorectification and image quality enhancement to standardize images for segmentation analysis. The pipeline was developed to autonomously execute the traditional GVF Snake algorithm, and strategically integrate SAM-P through automatically generated prompts directly targeting individual floes. Performance was quantitatively assessed using a benchmark dataset consisting of 20 manually segmented images representing over 2800 floes under varying lighting, weather, and ice conditions. The evaluation highlighted the relative strengths and weaknesses of each method by assessing their ability to identify and segment individual floes, estimate floe size distribution, and compute sea ice concentration. To our knowledge, this work presents the first evaluation of sea ice segmentation methods using a large-scale, manually annotated benchmark, and introduces an automated framework that enables reproducible, scalable, and robust segmentation analyses across diverse sea ice conditions.
@@
The GVF method was found to provide a reasonable floe size distribution due to its ability to capture the shape of floes, but was less effective at detecting all floes, impacting its accuracy in estimating ice concentration. SAM-A was found to perform well on images with distinct floe edges, but its performance deteriorates in more challenging environments. Throughout the sample, SAM-A detects the fewest floes, and its floe size distribution deviates most from the benchmark. Interestingly, it provides an accurate estimate of sea ice concentration, though this result is ambiguous as it arises from compensating factors: fewer floes are detected, but the identified floes are larger. In contrast, SAM-P is more effective at detecting floes and produces a more accurate size distribution. Its performance is consistent across the sample, resulting in a robust estimate of sea ice concentration. The SAM-GVF combination slightly improves SAM-P’s performance by reducing misinterpretation of complex floes, though its overall performance does not notably differ from SAM-P, returning size distributions and sea ice concentration that closely align with the benchmark.
@@
SAM-P and SAM-GVF emerged as the most effective methods, demonstrating strong performance across multiple metrics, particularly in accurately estimating sea ice concentration and improving floe detection, while maintaining good agreement with the benchmark in terms of floe size distribution. SAM-P is computationally fast, making it well-suited for real-time applications, where immediacy is critical, and for processing large imagery datasets to produce statistical properties of sea ice. In contrast, the combined SAM-GVF approach, along with GVF alone, proves more effective at capturing the specific floe size distribution. However, their longer computational time makes them more suitable for data refinement, where the specific information of the shape of floes is demanded.

### v4

@@ 4.3 A combination of SAM and GVF (SAM-GVF)
Our analysis indicates that the GVF algorithm efficiently delineates floes but encounters difficulties with varied gradients in gray-scale intensity due to its reliance on visual characteristics. SAM, on the other hand, can process a wide range of image variations due to its extended training, but it lacks shape integration, making it difficult to segment adjacent floes with weak edges. We propose combining SAM with GVF to compensate for shortcomings of the former by incorporating shape details, an approach conceptually similar to those suggested previously (e.g., Chen et al., 2019; Hoang Ngan Le et al., 2020).
@@ 5.1 Benchmark dataset
A benchmark is created by manually outlining a subset of twenty images from the original dataset (see §B). These images were sampled at random and then reviewed to ensure they excluded scenes dominated by open water or consolidated pack ice. The final set span a wide range of visual conditions and varying levels of sea ice clarity. Segmentation of individual floes is performed using image editing software (Pixelmator, though other tools could also be used). While this manual process provides high-quality reference data, it is inherently subject to a degree of subjectivity, especially in complex ice conditions where floe boundaries may be ambiguous. However, when performed by trained sea ice observers—in this study, three experienced observers who cross-checked a few sample images—the segmentation results were generally consistent and reproducible. Furthermore, to ensure temporal consistency in challenging cases, observers also referred to video recordings captured by the camera, cross-checking ambiguous floe boundaries against subsequent frames.
Reference data for the number and size of floes, as well as ice concentration, are subsequently derived. Concurrently, the same sample is segmented using the automated segmentation methods (GVF, SAM-A, SAM-P, and the combined SAM-GVF). Over 2,800 floes were identified across the 20 images, providing a solid foundation for benchmarking (Arbelàez et al., 2011; Martin et al., 2001).
@@ 5.2 Visual analysis of segmentation performance on sample images
Two sample images representing contrasting ice morphology, one with distinct ice floes and well-defined edges, and another featuring diffuse boundaries due to closely packed or interconnected floes and interstitial ice, are used to assess the different methods against manual segmentation (Figures 7 and 8). This comparison illustrates how varying sea ice conditions influence segmentation outcomes and provides insight into model performances (within their respective constraints). These include: missing floes, where floes identified in the manual segmentation are not detected by the automated methods; misinterpreted floes, which are either partially segmented or merged with adjacent floes; and “hallucinations”, referring to false objects where patterns or textures are mistakenly identified as floes (Tables 1 and 2).
@@ 5.2.1 Distinct ice floes with defined edges
SAM-GVF misses ten floes, misinterprets four, and produces seven hallucinations ($\approx$10% of floes). The missing floes have an average diameter of 2.6 m, covering around 1% of the surface area, similar to SAM-A and SAM-P. Misinterpreted floes have an average diameter of 3.4 m and include both partially segmented and merged floes. Hallucinations have an average diameter of 1.8 m and align with those seen in SAM-P.
Overall, SAM-GVF has the lowest percentage of segmentation errors, with uncertainties affecting <<3% of the image area. SAM-P achieves similar performance ($\approx$5%), while SAM-A and GVF exhibit the highest uncertainty, affecting 11% and 12% of the surface area, respectively.
@@ 5.2.2 Diffuse sea ice boundaries
As expected, segmentation performance declines under more challenging conditions. For the sample image with diffuse boundaries, the manual segmentation detects 250 floes with an average diameter of 3 m and an ice concentration of 55%. The performance of GVF significantly degrades—it misses 104 floes, nearly three times the number observed in the distinct edges image. Misinterpretations also increase, with 40 floes affected, while the 17 hallucinations is similar to the distinct edges case (Fig. 8g). In total, $\approx$70% of floes are not interpreted correctly, where the missed floes have an average diameter of 1.8 m and cover approximately 7% of the image area. The misinterpretations have an average diameter of 3 m and result mostly from partial segmentation of larger floes. The hallucinations have an average diameter of 1.1 m and do not substantially affect the overall sea ice characteristics. However, the increased number of missed and misinterpreted floes reduces the reliability of ice concentration estimates in the diffuse boundaries image.
@@
SAM-A is more robust for the diffuse boundaries image in comparison to GVF, although it shows a significant decline in performance. Approximately 81 floes are missed, which is over five times more than for the distinct edges image. Misinterpretations remain relatively stable at twenty floes, while hallucinations increase to ten, although this is still fewer than GVF (Fig. 7e). Overall, around 45% of floes are not captured correctly. The missed floes have an average diameter of 1.3 m, covering about 3% of the image area. While SAM-A misses more floes in the diffuse boundaries image compared to the distinct edges image, it still retains the capacity to provide an accurate estimate of ice concentration. Misinterpretations have an average floe diameter of 4.5 m, and primarily arise from the merging of adjacent floes, continuing SAM-A’s tendency to bias floe size distribution toward larger floes. Despite the increased hallucinations, their overall impact on sea ice characteristics remains limited.
@@
SAM-P also experiences an increase in segmentation errors, missing 53 floes, which is greater than six times more than in the distinct edges sample. The missed floes are small, with an average diameter of 1.2 m and covering only 2% of the total area. Misinterpretations double to 22 instances, involving smaller floes (average diameter of 3.3 m) than those in the distinct edges image. These misinterpretations primarily result from partial segmentation of floes featuring both snow-covered and bare ice surfaces, which subsequently affects the estimated floe size. Hallucinations increase slightly to 16, likely due to the model’s sensitivity to complex textures when using prompts to refine uncertain regions. Approximately 36% of floes are not interpreted correctly.
@@
SAM-GVF shows moderate performance degradation, missing 58 floes, which is a substantial increase (almost sixfold) compared to the distinct edges image. Misinterpretations remain relatively low, with 20 floes affected, while hallucinations increase slightly to 12, which is fewer than SAM-P and GVF (Fig. 8g). Similar to SAM-P, $\approx$36% of the floes are not captured correctly. The missed floes have an average diameter of 1.2 m, covering about 2% of the image area. Misinterpretations, with an average floe diameter of 2.8 m, mainly result from partial segmentation of larger floes and those with mixed snow and bare ice surfaces, which slightly distorts the floe size distribution. Hallucinations remain small, with an average diameter of 1.5 m, and do not significantly impact the overall sea ice characteristics.
@@
Of the various methods, SAM-P provides the best overall balance of accuracy and computational efficiency across both the distinct edges and diffuse boundaries images. SAM-P has the fewest missed floes and maintains stable, though slightly higher, rates of misinterpretations and hallucinations compared to SAM-GVF. While SAM-GVF performs marginally better in terms of misinterpretation and hallucination metrics, these improvements come at a substantial cost in computational time ($\approx$15 times slower per floe). SAM-A and GVF show significant deterioration, particularly with a substantial increase in missed floes. Additionally, computational time per floe increases for GVF (and consequently for SAM-GVF) in the diffuse boundaries image, whereas it remains steady in SAM. A comprehensive analysis performed on the entire benchmark dataset is presented in Table 4, confirming these findings.
@@ 5.3 Floe count, size distribution, and sea ice concentration across the entire sample
Data extracted from all 20 images in the sample are compared with the benchmark manual segmentation, as illustrated in Fig. 9. The relative error (RE) in estimating floe numbers is presented in Fig. 9a. Among the methods evaluated, SAM-P performs the best, slightly underestimating the number of floes by approximately 2%, which highlights its capability to detect floes across images with varying clarity. In contrast, SAM-A and GVF are the least accurate, underestimating the manual count by approximately 28% and 26%, respectively. This reflects the difficulty in detecting floes in complex scenarios, particularly when dealing with unclear or irregular ice formations (cf. Figs. 7 and 8). For SAM-A, the automated initialization is a key limitation, as it prioritizes larger, more distinct floes while often overlooking smaller or poorly defined ones. The combined SAM-GVF approach mitigates some of the uncertainties in floe detection observed in both SAM-A and GVF, leading to a lower underestimation of the manual count (approximately 12%). This aligns with its performance on the distinct edges image, suggesting that, like SAM-P, it remains relatively consistent across varying image conditions.
Figure 9: Comparison of floe size statistics derived from segmentation methods GVF (green), SAM-A (red), SAM-P (pink), and SAM-GVF (purple) against the manual benchmark (orange). (a) Relative error (RE) in the number of detected floes compared to the manual benchmark (horizontal orange line). b) Violin plots of floe size distributions for each method, with the manual benchmark on the left. Squares inside violins represent medians; darker shaded areas indicate quartiles. (c,d,e,f) $Q$ plots comparing deciles (10% to 90%) of floe size distributions from segmentation methods against the manual benchmark. The solid line indicates perfect alignment. (g) Relative error (RE) in ice concentration compared to the manual benchmark (horizontal orange line).
@@
The floe size distribution is shown in Fig. 9b as violin plots, which display the overall distribution of data, including the median (marked inside the violins) and the first and third quartiles (represented by the darker shaded areas). The benchmark size distribution from manual segmentation exhibits a clear bimodal shape, with peaks around diameters of 1.5 m and 3.5 m. Approximately 50% of the floes fall between 1.8 m and 4.3 m, with a median diameter of 3 m. SAM-P aligns well with the benchmark, with the bimodal shape, median, and quartiles generally well captured. However, the distribution is slightly skewed toward smaller floes, as the first quartile underestimates the benchmark (1.5 m instead of 1.8 m), exaggerating the presence of smaller floes and reflecting the relatively large count of hallucinations (Fig. 7g). Both GVF and SAM-GVF methods also capture the overall shape of the distribution, including the median (3.2 m and 3.1 m, respectively) and quartiles. Minor discrepancies arise from the lower detection of smaller floes, resulting in a less pronounced bimodal shape compared to SAM-P. Conversely, SAM-A produces the least accurate distribution, missing the bimodal nature of the benchmark, resulting in a single peak around 3.3 m. It overestimates both the median (3.8 m) and quartiles, reflecting the limitations of SAM-A in detecting individual floes without seed-based prompts.
@@
The comparison between benchmark and model distributions is further illustrated and analyzed in Figs. 9d-g, which present the deciles of the distributions as QQ plots. Despite differences in the overall bimodal shape, the deciles of SAM-P, SAM-GVF, and GVF align reasonably well with the benchmark (Figs. 9c,e,f). The lowest deciles (10–30%) are slightly more pronounced, especially for SAM-P, suggesting that these methods detect a large number of small floes, attributed to spurious features or noise. However, the higher deciles align well with the benchmark. The agreement between the deciles can be further quantified using the root mean squared error (RMSE) of floe diameter, calculated as the square root of the mean squared differences between the observed (manual benchmark) and predicted (segmentation method) deciles of the floe size distribution. Using this metric, the RMSE is 0.20 m for SAM-P, 0.17 m for SAM-GVF, and 0.21 m for GVF, which corresponds to approximately 3–4 times the smallest measurable unit in the images, given the 0.05 m pixel-to-meter conversion in the orthorectification process. SAM-A (Fig. 9d) shows a more significant underestimation compared to the other models, especially in the lowest five deciles. The RMSE for SAM-A is 0.94 m, reflecting its misalignment with the benchmark’s bimodality.
@@ 6 Conclusion
Three segmentation approaches have been evaluated for analyzing close-scale optical imagery of ice floes in the winter Antarctic MIZ. The first approach employs the gradient vector flow (GVF) snake algorithm, which uses edge-based energy minimization for contour detection. The second approach utilizes the Segment Anything Model (SAM), a deep learning framework for image segmentation, applied in both its automatic mode (SAM-A) and a prompt-guided approach (SAM-P). The third approach combines SAM-P with the GVF algorithm (SAM-GVF) to leverage their complementary strengths. To ensure robust and systematic evaluation, we implemented a fully automated image processing pipeline, incorporating orthorectification and image quality enhancement to standardize images for segmentation analysis. The pipeline autonomously executes the GVF snake algorithm and strategically integrates SAM-P through automatically generated prompts directly targeting individual floes. Performance was quantitatively assessed using a benchmark dataset consisting of 20 manually segmented images representing over 2800 floes under varying lighting, weather, and ice conditions. The evaluation highlighted the relative strengths and weaknesses of each method by assessing their ability to identify and segment individual floes, estimate floe size distribution, and compute sea ice concentration. To our knowledge, this work presents the first evaluation of sea ice segmentation methods using a large-scale, manually annotated benchmark, and introduces an automated framework that enables reproducible, scalable, and robust segmentation analyses across diverse sea ice conditions.
@@
The GVF method provided a reasonable floe size distribution due to its ability to capture floe shapes but was less effective in detecting all floes, impacting its accuracy in estimating ice concentration. SAM-A performed well on images with distinct floe edges, though its effectiveness decreased significantly in challenging environments, detecting fewer floes and providing the most divergent size distributions from the benchmark. SAM-A provided an accurate sea ice concentration estimate, albeit due to compensating factors: fewer, but larger detected floes. SAM-P was more effective at detecting floes, consistently producing accurate size distributions and robust sea ice concentration estimates across the entire sample. The SAM-GVF combination slightly improved upon SAM-P by reducing misinterpretations of complex floes, although overall performance remained similar, closely aligning with the benchmark in both floe size distribution and sea ice concentration.
@@
SAM-P and SAM-GVF emerged as the most effective methods, demonstrating strong performance across multiple metrics, especially in accurately estimating sea ice concentration and improving floe detection. SAM-P’s computational efficiency makes it suitable for real-time applications requiring immediate results and for processing large datasets to derive statistical properties of sea ice. Conversely, SAM-GVF and GVF alone, despite their longer computational times, are better suited for detailed analyses where precise floe shape information is essential.
This study establishes robust methodological foundations for obtaining reliable quantitative data on Antarctic sea ice floe sizes in the MIZ from field-acquired imagery, addressing the scarcity and limitations inherent in direct field observations. Although SAM, as a generalist segmentation model, is not currently optimized for sea ice, this presents an opportunity for further training specifically for Antarctic sea ice conditions using our extensive dataset. Building on the demonstrated effectiveness of these segmentation methods, future research will extend the analysis to include the complete image collection from numerous Antarctic expeditions and various seasons, enabling comprehensive monitoring of the evolution of Antarctic sea ice.